OpenStudy (priscilla):

how can i solve -1.5x+y=5 x+y=5

OpenStudy (anonymous):

I would multiply the entire first equation by (-1) so then your two equations would be 1.5x-y=-5 x+y=5 then you can use the combination method. add em all up: 2.5x=0 x=0 then if x=0 then you can plug it in 0+y=5 y=5 if you want to use the substitution method, you can solve one equation for either x or y, i would choose the second for neatness x+y=5 -x -x y=5-x then plug this new equation into the first one in lue of y 1.5x+x=5 2.5 x=5 divide, x still equals 0 then plug back in x and again you acheive y=5

OpenStudy (oaktree):

akingnh, she's been gone for 13 minutes now, so you might as well give up on that.

OpenStudy (anonymous):

she might come back! even so, try to help her out for future reference.